OpenStudy (anonymous):

From which graph can you best find the solution to the equation 7/x - 10/x^2 = 1

OpenStudy (campbell_st):

well multiply every term by x^2 and you get 7x - 10 = x^2 or x^2 - 7x + 10 = 0 hope it helps....

OpenStudy (anonymous):

kind of i need the graph for those